# On measures derived from orbital integrals

### v1

@@ 3. The projection of an orbital measure to the Cartan algebra
Let $\nu_{a}$ be the orbital measure on $\operatorname{\mathcal{O}}_{a}$ and define the measure $\mu_{a}$ on the Cartan algebra as $\mu_{a}=\operatorname{pr}_{*}(\nu_{a})$ where $\operatorname{pr}:\operatorname{\mathfrak{g}}\to\operatorname{\mathfrak{t}}$ is the orthogonal projection. The measure $\mu_{a}$ was characterized in [Z16, O13, F15], it can also be characterized as Duistermaat-Heckman measures in the context of Hamiltonian torus actions, see Section $5$ in the book [GLS96] and the references therein. We start with an equation satisfied by the measure $\mu_{a}$.
@@
For $a\in\operatorname{\mathfrak{t}}$ the measure $\mu_{a}$ satisfies
@@
Since $\langle\operatorname{Ad}_{g}x,y\rangle=\langle\operatorname{pr}(\operatorname{Ad}_{g}x),y\rangle$ for $y\in\operatorname{\mathfrak{t}}$ it is easy to very that the measure $\mu_{a}=\operatorname{pr}_{*}(\nu_{a})$ satisfies
@@
Therefore, using the orbital integral (1) the measure $\mu_{a}$ satisfies
@@
For $a\in\operatorname{\mathfrak{t}}$ the measure $\mu_{a}$ is characterized by
@@
We write $\eta=\left[\Delta,\Delta\right]/\Delta(a)$ and $l=|\Phi^{+}|$. If we take the part of degree not greater than $k\in\operatorname{\mathbb{N}}_{0}$ in $x$ of the equation (2) of the previous proposition then we get
@@
$\left[\int_{\operatorname{\mathfrak{t}}}q_{c}^{k}d\mu_{a}(c),f_{\operatorname{sym}}\right]$ $=\left[\int_{\operatorname{\mathfrak{t}}}q_{c}^{k}d\mu_{a}(c),f_{\operatorname{sym}}\right]$
$=\int_{\operatorname{\mathfrak{t}}}\left[q_{c}^{k},f_{\operatorname{sym}}\right]d\mu_{a}(c)$
@@
where we used Proposition 2.2 in the third equality and the symmetry of the measure $\mu_{a}$ in the final equality. Hence, the conclusion of the theorem holds. ∎
@@
Let us check that the measure $\mu_{a}$ characterized in this way is a probability measure. If we take the $f=\mathbf{1}$ then by Theorem 3.2 and Proposition 2.10
@@
$\int_{\operatorname{\mathfrak{t}}}f(c)d\mu_{a}(c)$ $=\eta\operatorname{ev}_{a}\left(I_{\Delta}f_{\operatorname{sym}}\right)$
$=\eta\operatorname{ev}_{a}\left(\int^{\langle x,\alpha^{\circ}\rangle}_{0}f_{\operatorname{sym}}(x-\alpha t)dt\right)$
$=\eta\int^{\left\langle a,\frac{1}{2}\alpha\right\rangle}_{0}f_{\operatorname{sym}}(a-\alpha t)dt$
@@
It is easy to see that the density of the measure $\mu_{a}$ is given by an indicator function in the interval $\left[a,-a\right]$ of $\operatorname{\mathfrak{t}}$. Since $\operatorname{\mathcal{O}}_{a}=\operatorname{Ad}_{\operatorname{SU}(2)}(a)\simeq S^{2}$ this is the fact known to Archimedes that the projection of the surface measure of a sphere to an axis is given by an indicator function.
@@ 4. The radial part of the convolution of orbital measures
Let $\nu_{a}$ and $\nu_{b}$ be the orbital measures on $\operatorname{\mathcal{O}}_{a}$ and $\operatorname{\mathcal{O}}_{b}$ for $a,b\in\operatorname{\mathfrak{t}}$. We characterize the radial part $\nu_{a,b}$ of the convolution of measures $\nu_{a}*\nu_{b}$ .It was characterized in previous research in [F19, IZ80, CMZ19] using mainly Fourier analytic techniques. We start with an equation satisfied by the radial measure $\nu_{a,b}$.
@@
For $a,b\in\operatorname{\mathfrak{t}}$ such that $\Delta(a)\Delta(b)\neq 0$ the measure $\nu_{a,b}$ satisfies
(4) $\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(b)\Delta(x)}$ $\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(a),x\rangle}\right)\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(b),x\rangle}\right)$
@@
If we express the Fourier-Laplace transform terms terms $\operatorname{\mathcal{F}}_{a},\operatorname{\mathcal{F}}_{b}$ and $\operatorname{\mathcal{F}}_{c}$ using the orbital integral formula (1), and if we multiply both sides by $\Delta(x)$ and divide both sides by $\left[\Delta,\Delta\right]$ then we obtain the equation (4) in the statement of the proposition. ∎
@@
(5) $\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}f(c)d\nu_{a,b}(c)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(b)}\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)T_{w(a)}P_{\operatorname{alt}}f\right)\right).$
for every polynomial $f\in\operatorname{\mathcal{P}}(\operatorname{\mathfrak{t}})$,
@@
$\varphi.F^{k}M_{r_{a}^{k}}\frac{r_{b}^{k+l}(x)}{\Delta(x)}=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}r_{c}^{k}(x)d\nu_{a,b}(c),$
If for a polynomial $f\in\operatorname{\mathcal{P}}^{k}(\operatorname{\mathfrak{t}})$ we take the inner product of both sides with $f_{\operatorname{alt}}=P_{\operatorname{alt}}f$ then on the LHS we have
@@
where we used the $r_{a}^{k}=P_{\operatorname{alt}}q_{a}^{k}$ in the first equality, Proposition 2.5 in the third, Proposition 2.9 in the fifth, Lemma 2.3 in the sixth and Proposition 2.2 in the final equality. On the other hand, on the RHS we get
@@
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}\left[r_{c}^{k}(x),f\right]d\nu_{a,b}(c)$
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}\operatorname{ev}_{c}(f)d\nu_{a,b}(c)$
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}f(c)d\nu_{a,b}(c),$
where we used Lema 2.3 in the second equality and Proposition 2.4 in the third. Combining the result of taking the inner product on both sides we obtain the equation in the statement of the theorem. ∎
Remark 4.3.
We see from equation (5) that the measure $\frac{1}{\Delta(c)}d\nu_{a,b}(c)$ is alternating. For a polynomial $f\in\operatorname{\mathcal{P}}(\operatorname{\mathfrak{t}})$ define $(w.f)(x)=f(w^{-1}(x))$ for $w\in\operatorname{\mathcal{W}}$. Since $P_{\operatorname{alt}}(w.f)=\epsilon(w)P_{\operatorname{alt}}(f)$ we have
$\int_{\operatorname{\mathfrak{t}}}(w.f)(c)\frac{1}{\Delta(c)}d\nu_{a,b}(c)=\epsilon(w)\int_{\operatorname{\mathfrak{t}}}f(c)\frac{1}{\Delta(c)}d\nu_{a,b}(c).$
@@
Consider the case in which $G=\operatorname{SU}(2)$, and $\alpha\in\operatorname{\mathfrak{t}}\subseteq\operatorname{\mathfrak{su}}(2)$ is given by $\alpha=e_{2}-e_{1}$ where $\{e_{1},e_{2}\}$ is the standard basis. Note that $\alpha^{\circ}=\frac{1}{2}\alpha$. For $f\in\operatorname{\mathcal{P}}(\operatorname{\mathfrak{t}})$
$\int_{\operatorname{\mathfrak{t}}}f(c)\frac{1}{\Delta(c)}d\nu_{a,b}(c)$ $=\varphi.\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{2}\left(T_{a}f_{\operatorname{alt}}-T_{-a}f_{\operatorname{alt}}\right)\right)\right)$
@@
$=\varphi\int^{\frac{1}{2}(b_{2}-b_{1})}_{0}\frac{1}{2}\left(f_{\operatorname{alt}}(b+a-\alpha t)-f_{\operatorname{alt}}(b-a-\alpha t)\right)dt$
@@
$\phi(c)=\begin{cases}\frac{c}{2ab}&\text{if $a-b\leq c\leq a+b$ or $-a-b\leq c\leq-a+b$}\\ 0&\text{otherwise.}\end{cases}$

### v2

@@ 3. The projection of an orbital measure to the Cartan algebra
For $a\in\operatorname{\mathfrak{t}}$ let $\nu_{a}$ be the orbital measure on $\operatorname{\mathcal{O}}_{a}$ and define the measure $\mu_{a}$ on the Cartan algebra as $\mu_{a}=\operatorname{pr}_{*}(\nu_{a})$ where $\operatorname{pr}:\operatorname{\mathfrak{g}}\to\operatorname{\mathfrak{t}}$ is the orthogonal projection. The measure $\mu_{a}$ was characterized in [Z16, O13, F15], it can also be characterized as a Duistermaat-Heckman measure in the context of Hamiltonian torus actions, see Section $5$ in the book [GLS96] and the references therein. We start with an equation satisfied by the measure $\mu_{a}$.
@@
For $a\in\operatorname{\mathfrak{t}}$ such that $\Delta(a)\neq 0$ the measure $\mu_{a}$ satisfies
@@
Since $\langle\operatorname{Ad}_{g}x,y\rangle=\langle\operatorname{pr}(\operatorname{Ad}_{g}x),y\rangle$ for $y\in\operatorname{\mathfrak{t}}$, $x\in\operatorname{\mathfrak{g}}$ and $g\in G$ it is easy to very that the measure $\mu_{a}=\operatorname{pr}_{*}(\nu_{a})$ satisfies
@@
Therefore, the orbital integral (1) implies that the measure $\mu_{a}$ satisfies
@@
For $a\in\operatorname{\mathfrak{t}}$ such that $\Delta(a)\neq 0$ the measure $\mu_{a}$ is characterized by
@@
We write $\eta=\left[\Delta,\Delta\right]/\Delta(a)$ and $l=|\Phi^{+}|$. If we take the part of degree not greater than $k\in\operatorname{\mathbb{N}}_{0}$ in $x$ of equation (2) of Proposition 3.1 then we have
@@
$\left[\int_{\operatorname{\mathfrak{t}}}q_{c}^{k}d\mu_{a}(c),f_{\operatorname{sym}}\right]$ $=\int_{\operatorname{\mathfrak{t}}}\left[q_{c}^{k},f_{\operatorname{sym}}\right]d\mu_{a}(c)$
@@
where we used Proposition 2.2 in the second equality and the symmetry of the measure $\mu_{a}$ in the final equality. Hence, the conclusion of the theorem holds. ∎
@@
Let us check that the measure $\mu_{a}$ characterized in this way is a probability measure. If we take as polynomial $f=\mathbf{1}$ then by Theorem 3.2 and Proposition 2.10
@@
$\int_{\operatorname{\mathfrak{t}}}f(c)d\mu_{a}(c)$ $=\eta\operatorname{ev}_{a}\left(I_{\Delta}P_{\operatorname{sym}}f\right)$
$=\eta\operatorname{ev}_{a}\left(\int^{\langle x,\alpha^{\circ}\rangle}_{0}P_{\operatorname{sym}}f(x-\alpha t)dt\right)$
$=\eta\int^{\left\langle a,\frac{1}{2}\alpha\right\rangle}_{0}P_{\operatorname{sym}}f(a-\alpha t)dt$
@@
It is easy to see that the density of the measure $\mu_{a}$ is given by the indicator function in the interval $\left[a,-a\right]$ of $\operatorname{\mathfrak{t}}$. Since adjoint orbits are spheres, that is, $\operatorname{\mathcal{O}}_{a}=\operatorname{Ad}_{\operatorname{SU}(2)}(a)\simeq S^{2}$ this is the fact known to Archimedes that the projection of the surface measure of a sphere to an axis is given by an indicator function.
@@ 4. The radial part of the convolution of orbital measures
Let $\nu_{a}$ and $\nu_{b}$ be the orbital measures on $\operatorname{\mathcal{O}}_{a}$ and $\operatorname{\mathcal{O}}_{b}$ for $a,b\in\operatorname{\mathfrak{t}}$. We characterize the radial part $\nu_{a,b}$ of the convolution of measures $\nu_{a}*\nu_{b}$. This was done in previous research in [F19, IZ80, CMZ19] using mainly Fourier analytic techniques. We start with an equation satisfied by the radial measure $\nu_{a,b}$.
@@
For $a,b\in\operatorname{\mathfrak{t}}$ such that $\Delta(a)\neq 0$ and $\Delta(b)\neq 0$ the radial measure $\nu_{a,b}$ satisfies
(4) $\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(b)}\frac{1}{\Delta(x)}$ $\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(a),x\rangle}\right)\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(b),x\rangle}\right)$
@@
If we express the Fourier-Laplace transform terms $\operatorname{\mathcal{F}}_{a},\operatorname{\mathcal{F}}_{b}$ and $\operatorname{\mathcal{F}}_{c}$ using the orbital integral formula (1), and if we multiply both sides by $\Delta(x)$ and divide both sides by $\left[\Delta,\Delta\right]$ then we obtain the equation (4) in the statement of the proposition. ∎
@@
For $a,b\in\operatorname{\mathfrak{t}}$ such that $\Delta(a)\neq 0$ and $\Delta(b)\neq 0$ the radial part $\nu_{a,b}$ of the convolution $\nu_{a}*\nu_{b}$ of orbital measures is characterized by
(5) $\int_{\operatorname{\mathfrak{t}}}\frac{f_{\operatorname{alt}}(c)}{\Delta(c)}d\nu_{a,b}(c)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(b)}\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)T_{w(a)}f_{\operatorname{alt}}\right)\right).$
for every alternating polynomial $f_{\operatorname{alt}}\in\operatorname{\mathcal{P}}_{alt}(\operatorname{\mathfrak{t}})$.
@@
$\varphi.F^{k}M_{r_{a}^{k}}\frac{r_{b}^{k+l}(x)}{\Delta(x)}=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}r_{c}^{k}(x)d\nu_{a,b}(c).$
For an alternating polynomial $f_{\operatorname{alt}}\in\operatorname{\mathcal{P}}_{\operatorname{alt}}^{k}(\operatorname{\mathfrak{t}})$ we take the inner product of both sides with $f_{\operatorname{alt}}$. On the LHS we have
@@
where we used the definition $r_{a}^{k}=P_{\operatorname{alt}}q_{a}^{k}$ in the first equality, Proposition 2.5 in the third, Proposition 2.9 in the fifth, Lemma 2.3 in the sixth and Proposition 2.2 in the final equality. On the other hand, on the RHS we get
@@
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}\operatorname{ev}_{c}(f_{\operatorname{alt}})d\nu_{a,b}(c)$
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}f_{\operatorname{alt}}(c)d\nu_{a,b}(c),$
where we used Proposition 2.4 in the second equality. Combining the result of taking the inner product on both sides we obtain the equation in the statement of the theorem. Since all symmetric polynomials are of the form $f_{\operatorname{alt}}/\Delta$ for an alternating polynomial $f_{\operatorname{alt}}$ and since the measure $\nu_{a,b}$ is symmetric we see that (5) completely characterizes the measure $\nu_{a,b}$. ∎
We can get a slight variation of the characterization in Theorem 4.2 of the radial measure $\nu_{a,b}$.
Corollary 4.3.
$\int_{\operatorname{\mathfrak{t}}}g(c)d\nu_{a,b}(c)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(b)}\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)T_{w(a)}(\Delta.P_{\operatorname{sym}}g)f\right)\right)$
for every polynomial $g\in\operatorname{\mathcal{P}}(\operatorname{\mathfrak{t}})$.
Proof.
An alternating polynomial $f_{\operatorname{alt}}\in\operatorname{\mathcal{P}}_{\operatorname{alt}}(\operatorname{\mathfrak{t}})$ can be written as $f_{\operatorname{alt}}=\Delta.P_{\operatorname{sym}}g$ for a polynomial $g\in\operatorname{\mathcal{P}}(\operatorname{\mathfrak{t}})$. Making this substitution in formula (5) and noting that
$\int_{\operatorname{\mathfrak{t}}}(P_{\operatorname{sym}}h)(c)d\nu_{a,b}(c)=\int_{\operatorname{\mathfrak{t}}}h(c)d\nu_{a,b}(c)$
for any polynomial $h\in\operatorname{\mathcal{P}}(\operatorname{\mathfrak{t}})$ the corollary follows. ∎
@@
Consider the case in which $G=\operatorname{SU}(2)$, and $\alpha\in\operatorname{\mathfrak{t}}\subseteq\operatorname{\mathfrak{su}}(2)$ is given by $\alpha=e_{2}-e_{1}$ where $\{e_{1},e_{2}\}$ is the standard basis. Note that $\alpha^{\circ}=\frac{1}{2}\alpha$. For $f_{\operatorname{alt}}\in\operatorname{\mathcal{P}}_{\operatorname{alt}}(\operatorname{\mathfrak{t}})$
$\int_{\operatorname{\mathfrak{t}}}f_{\operatorname{alt}}(c)\frac{1}{\Delta(c)}d\nu_{a,b}(c)$ $=\varphi.\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{2}\left(T_{a}f_{\operatorname{alt}}-T_{-a}f_{\operatorname{alt}}\right)\right)\right)$
@@
$=\varphi.\int^{\frac{1}{2}(b_{2}-b_{1})}_{0}\frac{1}{2}\left(f_{\operatorname{alt}}(b+a-\alpha t)-f_{\operatorname{alt}}(b-a-\alpha t)\right)dt$
@@
$\phi(c)=\begin{cases}\frac{c}{4ab}&\text{if $a-b\leq c\leq a+b$}\\ \frac{-c}{4ab}&\text{if $-a-b\leq c\leq-a+b$}\\ 0&\text{otherwise.}\end{cases}$